OpenStudy (anonymous):

If 7p + 5q = 2 and 8p - 9q = 17, then -p + 14q =

OpenStudy (solomonzelman):

First you have a system of equations. 7p + 5q = 2 8p - 9q = 17 What methods to solve those do you know?

OpenStudy (anonymous):

opposite coefficients?

OpenStudy (solomonzelman):

That means that you would do what?

OpenStudy (anonymous):

multiply the first equation by 9 and second one by 5

OpenStudy (solomonzelman):

Good!

OpenStudy (anonymous):

i dont get the last part

OpenStudy (solomonzelman):

OK, multiply the equations, 1st times 9 and second times 5. Tell me what you get.

OpenStudy (solomonzelman):

Do you need help with multiplying the equations?

OpenStudy (anonymous):

ok hold up

OpenStudy (anonymous):

63p=18

OpenStudy (solomonzelman):

I don't think so, do it again. (or you want me to show your mistake?) 7p + 5q = 2 8p - 9q = 17

OpenStudy (anonymous):

times the whole equation by 9 the Q cancels out and 7*9 is 63 and 9*2=18

OpenStudy (solomonzelman):

Yes and you are adding the result of the second equation to that. Don't just cancel the Q s.

OpenStudy (anonymous):

is it 1

OpenStudy (solomonzelman):

the p? How do you know that?

OpenStudy (anonymous):

i solved it

OpenStudy (solomonzelman):

Cool@!

OpenStudy (solomonzelman):

Ok, you got 103p=103 right?

OpenStudy (anonymous):

Shalom brother

OpenStudy (solomonzelman):

And you can do the second variable. DON"T forget that you are SOLVING for -p + 14q, OK?

OpenStudy (solomonzelman):

Shalom, how do you know I am Jewish?

OpenStudy (anonymous):

gotcha!

OpenStudy (solomonzelman):

U did get me!

OpenStudy (anonymous):

lol a brother recognizes a smart brother

OpenStudy (solomonzelman):

Aleychem Shalom, then!

OpenStudy (solomonzelman):

You are a 4th Jew I met in the OS.

OpenStudy (anonymous):

nice

OpenStudy (solomonzelman):

להתראות

OpenStudy (solomonzelman):

I met Asher, (that one is obvious) you and some other girl with a nickname. I am going to fan you.

OpenStudy (anonymous):

ok go ahead

OpenStudy (solomonzelman):

Are you sephardy or ashkenaz?

OpenStudy (solomonzelman):

sefardy, my bad.